OpenStudy (anonymous):

this is a pretty complicated question if someone could explain it i would appreciate it

OpenStudy (phi):

Does this have a figure?

OpenStudy (anonymous):

its a cone and a sphere

OpenStudy (anonymous):

im getting 4pisqrt(2) but thats wrong

OpenStudy (anonymous):

fx is (x)*sqrt(x^2+y^2) and fy is (x)*sqrt(x^2+y^2)

OpenStudy (phi):

yes that is what I got, for the outer surface of the sphere. Do they want to include the surface area of the inside of the sphere?

OpenStudy (anonymous):

but the answer is wrong the correct answer is4π(2−2sqrt(2))

OpenStudy (phi):

just to make sure I am interpreting the problem correctly, do they have a figure?

OpenStudy (anonymous):

no they don't =(

OpenStudy (anonymous):

the equation of a sphere is z^2+x^2+y^2=4 does this have anything to do with this prob??

OpenStudy (phi):

I would use polar coordinates... but I'll puzzle on this, and post something later.

OpenStudy (anonymous):

alright thanks=)

OpenStudy (phi):

First, you say the correct answer is4π(2−2sqrt(2)) that gives a negative number But if you use the idea that the surface of the sphere is the area of a set of "rings" with circumference r and surface area ds

OpenStudy (anonymous):

it is

OpenStudy (phi):

 if we integrate from 0 to pi we get the surface area of a sphere, 4 pi r^2 Here we integrate from 0 to pi/4 where the sphere meets the cone we get \[ 2 \pi r^2 (1 - \frac{\sqrt{2}}{2}) \] with r= 2 this is \[ 8 \pi (1 - \sqrt{2}/2) \] or \[ 4 \pi (2 - \sqrt{2}) \]